### v2

# GLUE: Gradient-free Learning to Unify Experts

Abstract
In many deployed systems (multilingual ASR, cross-hospital imaging, region-specific perception), multiple pretrained specialist models coexist. Yet, new target domains often require domain expansion: a generalized model that performs well beyond any single specialist’s domain. Given a new target domain, existing methods obtain a single strong initialization prior for the model parameters by blending expert models to initialize a target model. However, heuristic blending—using mixing coefficients based on data size or proxy metrics—often yields lower target-domain test accuracy, and learning these coefficients on the target domain’s loss function typically requires computationally-expensive full backpropagation through a neural network. We propose GLUE, Gradient-free Learning to Unify Experts, which initializes the target model as a convex combination of fixed experts and learns the mixture coefficients of this combination via gradient-free two-point SPSA (simultaneous perturbation stochastic approximation) updates, requiring only two forward passes per step. Across experiments on three datasets and three network architectures, GLUE produces model parameter priors that can be fine-tuned to outperform baselines. GLUE improves test accuracy by up to 8.5% over data-size weighting and by up to 9.1% over proxy-metric selection. GLUE either outperforms backpropagation-based full-gradient mixing or matches its performance within 1.4%.
Index Terms—  Expert Weight Mixing, Model Soup, Gradient-Free Optimization, Two-Point SPSA, Transfer Learning
1 Introduction
Pretrained parameters, whether from specialist “expert” models or large foundation models, enable practitioners to quickly start learning new tasks by using the pretrained parameters to initialize a target network and finetune it on the new task data [22, 38, 24]. This practice is common in multilingual or multi-accent automatic speech recognition (ASR) [32], radiology models deployed across scanners and hospitals [1], perception stacks that face region/weather variation [30, 23], recommendation systems partitioned by locale or user segments [7], and foundation model adapters specialized for code, legal, or biomedical text [40, 43, 24]. In these settings, initializing a target network with expert parameters often outperforms knowledge distillation [11, 4], which requires extra hyperparameters and can introduce optimization noise (e.g., temperature tuning, teacher-student mismatch). In contrast, directly initializing a target network with experts’ parameters simplifies downstream fine-tuning while preserving pretrained feature hierarchy and inductive biases [41, 17, 18, 19].
In general, adopting a single expert’s parameters is often suboptimal when building a general-purpose model, because real-world data are fragmented across related but distinct distributions [6, 29, 35]. Different experts capture complementary structures [39, 34, 3], motivating a parameter-space fusion of these experts as an initialization prior. Previous works [16, 25, 26, 5, 33, 28] form convex combinations of the $P$-dimensional parameters of $K$ expert models $\{\boldsymbol{\theta}_{i}\}_{i=1}^{K}$
$\boldsymbol{\theta}(\boldsymbol{\alpha})=\sum_{i=1}^{K}\alpha_{i}\boldsymbol{\theta}_{i},\;\boldsymbol{\theta}_{i}\in\mathbb{R}^{P}$ (1)
where the mixture coefficients ${\boldsymbol{\alpha}}=(\alpha_{1},\ldots,\alpha_{K})$ define a single deployable model tailored to the target domain.
Common heuristics for finding $\boldsymbol{\alpha}$ include (i) weighting experts by the size of their pretraining data [16], (ii) weighting by an external metric such as held-out accuracy on a proxy dataset[33, 15, 24], or (iii) learning the weights by optimizing the target loss[12, 36, 37]. Among these, learning the weights $\boldsymbol{\alpha}$ generally leads to the strongest single initialization (prior) for fine-tuning, because it optimizes the true target loss and can capture interactions among experts that global heuristics miss [20, 33].
The practical challenge is computational cost. Even when the number of mixture weights $K$ is small, standard gradient descent on the mixture coefficients still requires computing the loss gradient $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$ for the full parameter vector $\boldsymbol{\theta}(\boldsymbol{\alpha})$ of dimension $P\gg K$, thus requiring a complete forward and backward pass over the new model.
We propose Gradient-free Learning to Unify Experts (GLUE), a method for learning expert mixing coefficients $\boldsymbol{\alpha}$ via two-point simultaneous perturbation stochastic approximation (SPSA) [27, 10]. Starting from (1), GLUE optimizes only the low-dimensional vector $\boldsymbol{\alpha}$ [13, 14]. To do so, at each iteration, it samples a random direction $\mathbf{u}\in\mathbb{R}^{K}$, evaluates the loss at $\boldsymbol{\alpha}\pm\delta\mathbf{u}$ (two forward passes with blended weights $\boldsymbol{\theta}(\boldsymbol{\alpha}\pm\delta\mathbf{u})$), and updates $\boldsymbol{\alpha}$ using the finite-difference estimate. No backpropagation is required—only forward passes—making GLUE lightweight and easy to deploy. While SPSA has been used in other learning tasks, it may be inaccurate in high dimensions [27, 21, 10]. The key insight behind GLUE is that the mixture coefficients are only $K$-dimensional, mitigating this common drawback of SPSA. Our contributions are threefold:
[…]
In the following, we review related work (Section 2); introduce GLUE and its optimization procedure (Section 3); develop computational and statistical analyses (Section 4); present experiments against baselines (Section 5); discuss practical considerations and limitations (Section 6); and conclude (Section 7).
2 Related Work
A straightforward way to combine experts is to weight each model in proportion to the amount of data used to train it—an idea resembling that of federated averaging (FedAvg), where model updates from multiple clients are aggregated based on clients’ local dataset sizes in order to produce a unified model that (on average) performs well on all clients’ data [16]. While simple and scalable, this strategy is agnostic to the target domain. Under distribution changes, such as domain expansion, data-size weighting cannot adapt to target loss signals and can overemphasize large but misaligned sources.
Another common approach involves choosing or reweighting experts using a proxy metric (e.g., held-out accuracy) and then averaging the expert model parameters. Examples include “model soups”—uniform averaging or greedy selection before averaging—and Fisher-weighted merging, which uses per-parameter Fisher information with model-level mixing as a hyperparameter  [33, 15, 24]. These methods can be effective when a reliable proxy set exists, but they decouple weight choice from the training objective and often require manual tuning or discrete selection steps.
A third line of work learns the expert mixture coefficients by minimizing a target-driven objective, ranging from scalar task-vector combinations (TIES-Merging [36]) to interference-aware merging and adaptive (un)supervised schemes that tune model-wise coefficients (AdaMerging) [12, 37]. These gradient-based methods require computing $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$ via backpropagation through the blended network and then taking inner products with expert parameters (e.g., $\partial\mathcal{L}/\partial\boldsymbol{\alpha}_{i}=\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},\,\boldsymbol{\theta}_{i}\rangle$), which is computationally expensive.
Zeroth-order methods replace backpropagation with finite differences. To compute the two-point simultaneous perturbation stochastic approximation (SPSA) estimator [31], $\boldsymbol{\alpha}$ can be perturbed along a random direction to estimate a directional derivative with two forward passes, staying in the $K$-dimensional mixture space (where $K$ is the number of experts). However, the estimator’s variance grows with dimension and curvature, demanding many directions or delicate tuning and often causing slow or unstable convergence for large $K$ or ill-conditioned losses [42].
3 Methodology
In this section, we introduce GLUE, Gradient-free Learning to Unify Experts, which builds a single prior by mixing fixed experts and learning the mixing parameters $\boldsymbol{\alpha}$ via two-point SPSA.
3.1 Objective
Given $K$ pre-trained expert models with parameters $\{\boldsymbol{\theta}_{i}\}_{i=1}^{K}$, we aim to synthesize a single pretrained model $\boldsymbol{\theta}(\boldsymbol{\alpha})$ (defined in (1)) that serves as a strong prior for a new target domain that is related to the sources but exhibits a shifted data distribution. Throughout, we only optimize the low-dimensional mixing coefficients $\boldsymbol{\alpha}$; the expert parameters $\{\boldsymbol{\theta}_{i}\}_{i=1}^{K}$ remain fixed.
Let $f(\cdot;{\boldsymbol{\theta}})$ denote the model (neural network) instantiated with parameters $\boldsymbol{\theta}$. For a minibatch $\mathcal{B}=\{(\mathbf{x}_{j},\mathbf{y}_{j})\}_{j=1}^{B}$ from the target domain and a standard supervised loss $\ell$ (cross-entropy for classification or squared error for regression), we minimize empirical risk:
$\mathcal{L}(\boldsymbol{\alpha})=\frac{1}{B}\sum_{j=1}^{B}\ell\left(f(\mathbf{x}_{j};{\boldsymbol{\theta}(\boldsymbol{\alpha})}),\,\mathbf{y}_{j}\right)$
We treat $\boldsymbol{\alpha}$} as the sole optimization variable, and employ a zeroth-order (two-point) update on $\boldsymbol{\alpha}$.
3.2 Two-point SPSA updates
Since our optimization variable is the $K$-dimensional mixture, and zeroth-order estimates are accurate in lower dimensions [27, 21, 10], we estimate the gradient of $\mathcal{L}$ with respect to $\boldsymbol{\alpha}$ by using a two-point simultaneous perturbation stochastic approximator (SPSA).
At each iteration, we draw a random direction vector $\mathbf{u}\in\mathbb{R}^{K}$ (e.g., by drawing entries of $\mathbf{u}$ i.i.d. from a standard Gaussian distribution and normalizing) and choose a small radius $\mu>0$. We form perturbed unconstrained coefficients $\boldsymbol{\alpha}_{\pm}=\boldsymbol{\alpha}\pm\mu\mathbf{u}$ and corresponding blended parameters $\boldsymbol{\theta}(\boldsymbol{\alpha}_{\pm})=\sum_{i=1}^{K}{\alpha}_{{\pm}_{i}}\,\boldsymbol{\theta}_{i}$. At both blends, we then evaluate the loss on the same minibatch of target-domain examples, keeping inference deterministic across the two probes (e.g., BatchNorm in evaluation mode and fixed randomness) so that differences reflect only the perturbation:
$L_{\pm}=\frac{1}{b}\sum_{j=1}^{b}\ell\left(f(\mathbf{x}_{j};\boldsymbol{\theta}(\boldsymbol{\alpha}_{\pm})),\,\mathbf{y}_{j}\right)$ (2)
The directional finite difference along $\mathbf{u}$ is
$d(\boldsymbol{\alpha};\mathbf{u})=\frac{L_{+}-L_{-}}{2\mu}$
which induces the Gaussian-smoothing gradient estimate
$\widehat{\nabla}_{\boldsymbol{\alpha}}\mathcal{L}(\boldsymbol{\alpha})=d(\boldsymbol{\alpha};\mathbf{u})\,\mathbf{u}$
If desired, we can reduce variance by additionally averaging over $m$ independently sampled directions $\{\mathbf{u}_{r}\}_{r=1}^{m}$ in the same iteration,
$\widehat{\nabla}_{\boldsymbol{\alpha}}\mathcal{L}(\boldsymbol{\alpha})=\frac{1}{m}\sum_{r=1}^{m}\frac{L^{(r)}_{+}-L^{(r)}_{-}}{2\mu}\,\mathbf{u}_{r}$ (3)
with $\mathbf{u}_{r}$ sampled as above and $L^{(r)}_{\pm}$ computed as in (2). Finally, we update the mixture parameters:
$\boldsymbol{\alpha}\leftarrow\boldsymbol{\alpha}-\eta\widehat{\nabla}_{\boldsymbol{\alpha}}\mathcal{L}(\boldsymbol{\alpha})$ (4)
where $\eta>0$ is the step size. In our experiments, we perform SPSA and momentum updates in an unconstrained $\boldsymbol{\beta}$-space and map to $\boldsymbol{\alpha}=\mathrm{softmax}(\boldsymbol{\beta})$, replacing (4). Optimizing with respect to $\boldsymbol{\beta}$ ensures that $\boldsymbol{\theta}(\boldsymbol{\alpha})$ remains within the convex hull of expert parameters.
3.3 From mixture to prior
We initialize $\boldsymbol{\alpha}=\frac{1}{K}\mathbf{1}$ and iterate the two-point updates over minibatches until validation performance plateaus. Formally, we target the empirical-risk minimizer and map it to one blended model:
$\boldsymbol{\alpha}^{\star}\in\operatorname*{argmin}_{\boldsymbol{\alpha}\in\mathbb{R}^{K}}\mathcal{L}(\boldsymbol{\alpha}),\;\qquad\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}(\boldsymbol{\alpha}^{\star})$
The resulting $\boldsymbol{\theta}^{\star}$ serves as the starting point for standard gradient-based adaptation on the target domain, providing a strong prior that concentrates the experts’ most relevant inductive biases.
4 Theoretical Analysis
In this section, we show that the two-point (SPSA) procedure used in GLUE is (i) computationally cheaper per iteration than gradient-based mixing and (ii) statistically stable when the mixture dimension $K$ (the number of experts) is sufficiently small. To do so, we leverage results in stochastic approximation and zeroth-order optimization [27, 21, 10]; under standard smoothness and bounded-noise assumptions, two-point SPSA converges to first-order stationary points, so GLUE inherits these classical guarantees [8].
4.1 Cost for updating $\boldsymbol{\alpha}$}: two-point (SPSA) vs. full-gradient
We compare a single SPSA step on $\boldsymbol{\alpha}$} with a full-gradient step on $\boldsymbol{\alpha}$} when the model parameters are blended as ${\boldsymbol{\theta}}({\boldsymbol{\alpha}})=\sum_{i=1}^{K}\alpha_{i}{\boldsymbol{\theta}}_{i}$ and the experts $\{{\boldsymbol{\theta}}_{i}\}_{i=1}^{K}$ are fixed.
Let $F$ denote the cost of one forward pass of the blended network on a minibatch of data, and let the cost of a full backward pass be $\gamma F$ with $\gamma\geq 2$ (typical for deep neural networks [9]). Let $C_{\text{mix}}$ be the cost of the multiply-accumulate operations used to form a blended parameter vector ${\boldsymbol{\theta}}({\boldsymbol{\alpha}})$ from the $K$ experts: $C_{\text{mix}}=O(PK)$ memory operations for parameter dimension $P$. Let $D_{{\boldsymbol{\alpha}}}$ be the cost of computing the gradient $\nabla_{\boldsymbol{\alpha}}\mathcal{L}=\big[\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},{\boldsymbol{\theta}}_{1}\rangle,\ldots,\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},{\boldsymbol{\theta}}_{K}\rangle\big]$ given $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$, which is an $O(PK)$ dot-product accumulation. Evaluating $\nabla_{\boldsymbol{\alpha}}\mathcal{L}$ at the current mixture requires (i) computing ${\boldsymbol{\theta}}({\boldsymbol{\alpha}})$ once, (ii) one forward pass and one backward pass to compute $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$, and (iii) $D_{\boldsymbol{\alpha}}$ for computing $K$ inner products with $\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},\boldsymbol{\theta}_{i}\rangle$. With one perturbation pair ($m=1$), SPSA evaluates the loss at two nearby mixtures ${\boldsymbol{\alpha}}_{\pm}={\boldsymbol{\alpha}}\pm\delta\mathbf{u}$, requiring two blends and two forwards, but no backward. Therefore:
$T_{\text{full}}$ $=(1+\gamma)F+C_{\text{mix}}+D_{{\boldsymbol{\alpha}}},\qquad T_{\text{SPSA}}=2(F+C_{\text{mix}})$
Whenever $T_{\text{SPSA}}<T_{\text{full}}$, SPSA is cheaper than the full-gradient step (i.e., $C_{\text{mix}}<(\gamma-1)F+D_{\boldsymbol{\alpha}}$). In practice, the condition is easily met since (i) the inner product cost $D_{\boldsymbol{\alpha}}$ is typically of the same order as the mixing cost $C_{\text{mix}}$ and (ii) the cost of the backward pass is typically more than $2F$ ($\gamma\geq 2$).
4.2 Variance of two-point estimates
We analyze the variance of the two-point (SPSA) gradient estimator $\overline{\mathbf{g}}(\boldsymbol{\alpha})$, an average over $m$ i.i.d. random direction vectors $\{\mathbf{u}_{r}\}_{r=1}^{m}$:
$\overline{\mathbf{g}}(\boldsymbol{\alpha})$ $=\frac{1}{m}\sum_{r=1}^{m}\widehat{\mathbf{g}}(\boldsymbol{\alpha};\mathbf{u}_{r})$ (5)
$\widehat{\mathbf{g}}(\boldsymbol{\alpha};\mathbf{u})$ $=\left(\frac{{\mathcal{L}}(\boldsymbol{\alpha}+\mu\mathbf{u})-{\mathcal{L}}(\boldsymbol{\alpha}-\mu\mathbf{u})}{2\mu}\right)\mathbf{u},\;\;\mu>0$
Let $\mathbf{g}=\nabla_{\boldsymbol{\alpha}}\mathcal{L}$ denote the true gradient of the loss with respect to $\boldsymbol{\alpha}$. As $\mu\rightarrow 0$, $\widehat{\mathbf{g}}$ becomes the directional derivative of $\mathcal{L}$ with respect to $\mathbf{u}$ and the approximation error scales with the square of $\mu$: $\widehat{\mathbf{g}}=(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})$. Since the directional derivative is an unbiased estimator of $g$, both $\widehat{\mathbf{g}}$ and $\overline{\mathbf{g}}$ are also unbiased [2]. Therefore, we bound the variance of $\overline{\mathbf{g}}$ in the following proposition—where the variance refers to the trace of the covariance matrix of the averaged gradient estimator $\overline{\mathbf{g}}$.
Proposition 1 (Variance Bound).
The variance of the two-point (SPSA) gradient estimator $\overline{\mathbf{g}}(\boldsymbol{\alpha})$ is upper bounded as:
$\mathbb{E}\left[\|\overline{\mathbf{g}}-\mathbf{g}\|^{2}\right]$ $\leq\frac{K-1}{mK}\sigma_{\mathrm{max}}(\boldsymbol{\Theta})^{2}\|\nabla_{\boldsymbol{\theta}}\mathcal{L}\|^{2}+O\left(\mu^{2}\right)$
where $\boldsymbol{\Theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\dots,\boldsymbol{\theta}_{K}]$ and $\boldsymbol{\theta}=\boldsymbol{\Theta}\boldsymbol{\alpha}$.
Proof.
We consider the first order Taylor approximation of (5):
$\widehat{\mathbf{g}}(\boldsymbol{\alpha};\mathbf{u})$ $=(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})$
We proceed to compute $\mathbb{E}\left[\|\widehat{\mathbf{g}}-\mathbf{g}\|^{2}\right]$ by using the fact that $\mathbf{u}$ being a unit isotropic random vector implies $\|\mathbf{u}\|=1$ and $\mathbb{E}[\mathbf{u}\mathbf{u}^{\top}]=\frac{1}{K}\mathbf{I}$.
$\mathbb{E}$ $\left[\|\widehat{\mathbf{g}}-\mathbf{g}\|^{2}\right]=\mathbb{E}\left[\left\|(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})-\mathbf{g}\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}\right\|^{2}-2(\mathbf{g}^{\top}\mathbf{u})^{2}\right]+\|\mathbf{g}\|^{2}+O(\mu^{2})$
$=-\mathbf{g}^{\top}\mathbb{E}\left[\mathbf{u}\mathbf{u}^{\top}\right]\mathbf{g}+\|\mathbf{g}\|^{2}+O(\mu^{2})=\frac{K-1}{K}\|\mathbf{g}\|^{2}+O(\mu^{2})$
The variance with respect to $\overline{\mathbf{g}}$ is then:
$\mathrm{Var}[\overline{\mathbf{g}}]$ $=\frac{1}{m}\mathrm{Var}[\widehat{\mathbf{g}}]=\frac{K-1}{mK}\|\mathbf{g}\|^{2}+O(\mu^{2})$ (6)
We proceed to bound $\|\mathbf{g}\|^{2}$ as:
$\|\mathbf{g}\|^{2}=\|\nabla_{\boldsymbol{\alpha}}\mathcal{L}\|^{2}=\|\boldsymbol{\Theta}\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}\|^{2}\leq\sigma_{\mathrm{\max}}(\boldsymbol{\Theta})^{2}\|\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}\|^{2}$
Substituting the inequality above into (6) completes the proof. ∎
The proposition demonstrates that if $\sigma_{\mathrm{\max}}(\boldsymbol{\Theta})$ is small, e.g., the expert parameters are more aligned, the variance grows slower in the number of experts $K$. The variance can be reduced with additional random direction samples (increasing $m$). Interestingly, we find that $m=1$ already provides stable updates for small $K$ empirically (see Section 5), which is consistent with the $1/m$ scaling.
Additionally, assuming that the loss function $\mathcal{L}$ is Lipschitz smooth with respect to $\theta$ leads to an upper bound on $\|\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})\|$.
Fig. 1: Fine-tuning test accuracy vs. lightweight finetuning epochs (with fixed $\boldsymbol{\alpha}$). Datasets: (a) CIFAR-10, (b) SVHN, (c) Imagenette. We compare Configs 1-4 (see Sec. 5). GLUE (Config 4; ours) shows similar trends to Config 3 (full backpropagation for gradient computation) and consistently achieves higher test accuracy and faster convergence than Configs 1-2 (heuristics for $\boldsymbol{\alpha}$ based on data size and proxy accuracy).
5 Experimental Evaluations
We evaluate GLUE on CIFAR-10, SVHN, and Imagenette datasets using three backbones: ResNet-20, MobileNetV2, and an 8-layer ViT (patch size 8), respectively. We train $K=10$ expert models on heterogeneous splits of the base dataset. We adopt a non-IID experimental evaluation setup commonly used for federated learning: we sample per-class proportions for each expert from a Dirichlet distribution with concentration parameter $\delta=0.5$, then subsample a total of 2000 images to match those class proportions. In general, $\delta\leq 0.5$ induces skewed (non-IID) splits; larger $\delta$ approaches IID. Each expert is trained for 40 epochs with a batch size of 64 and an Adam optimizer with a learning rate of 0.001 and momentum parameters $(0.9,0.999)$. After training the experts, we evaluate four methods for determining $\boldsymbol{\alpha}$, the mixture coefficients used to combine the expert models’ parameters:
[…]
Configs 1 and 2 do not optimize on the target domain; Configs 3 and 4 learn $\boldsymbol{\alpha}$ using target-domain training data. In Configs 3 and 4, we use Adam optimizer with learning rate 1e-2, momentum parameters (0.9, 0.99) to update $\boldsymbol{\alpha}$. Throughout training, $\boldsymbol{\theta}(\boldsymbol{\alpha})$ is constrained to lie in the convex hull of experts (i.e., $\sum_{i}\alpha_{i}=1$). Additionally for GLUE, we update $\boldsymbol{\alpha}$ by randomly sampling only one random direction $u$. We use 10,000 images sampled IID from the base dataset to learn the unification parameters $\boldsymbol{\alpha}$ and finetune the models as described below. All unified models are subsequently evaluated on the original testing split of each of the base datasets.
After determining $\boldsymbol{\alpha}$, we form the unified initialization $\boldsymbol{\theta}(\boldsymbol{\alpha})$ for the target model and report: (i) prior (zero-shot) performance of $\boldsymbol{\theta}(\boldsymbol{\alpha})$ without any target training, and (ii) finetuned performance after updating all network parameters on the target training set while keeping $\boldsymbol{\alpha}$ fixed. We also report the test accuracy of the individual experts before blending in Table 1 for reference. We ran all finetuning experiments with three different random seeds and report the averaged test accuracy.
Table 1: Test accuracies (%) of the 10 experts evaluated on the target training dataset. We bolded the highest/lowest accuracies per row.
Figure 1 plots test accuracy during fine-tuning for the four configurations. Across all datasets and backbones, GLUE (Config 4) consistently outperforms data-size weighting (Config 1) by up to 6.7% on CIFAR-10, 3.8% on SVHN, and 8.5% on Imagenette. GLUE also outperforms proxy-accuracy weighting (Config 2) by up to 7.0% on CIFAR-10, 3.9% on SVHN, and 9.1% on Imagenette. GLUE attains similar performance to full-grad $\boldsymbol{\alpha}$ (Config 3): GLUE outperforms full-grad on CIFAR-10 by up to 4.5% and has at most 1.4% and 0.5% variation in test accuracy on SVHN and Imagenette, respectively. To minimize computational cost in these experiments, GLUE computes only one two-point estimate to approximate the gradient at each iteration ($m=1$). However, the variation in test accuracy between GLUE and full-grad $\boldsymbol{\alpha}$ (Config 3) can be reduced by averaging multiple two-point estimates to obtain better approximations to the true gradient—as in (3).
In summary, Figure 1 shows that GLUE (Config 4) starts from a stronger prior and converges faster to a higher test accuracy than the alternatives, closely tracking the performance of the full-gradient method (Config 3).
These results highlight that 1) the performance gains from learning $\boldsymbol{\alpha}$ persist even after finetuning the target model and 2) GLUE mimics the gains of full-gradient $\boldsymbol{\alpha}$ learning while providing computational benefits by relying only on forward evaluations instead of costly backpropagation.
6 Discussion
In GLUE, each step forms two blends and runs two forward passes; there is no backward pass. The per-step cost advantage over full-gradient mixing follows from
$T_{\text{full}}-T_{\text{SPSA}}$ $=(1+\gamma)F+C_{\text{mix}}+D_{\boldsymbol{\alpha}}-2(F+C_{\text{mix}})$
$=(\gamma-1)F-C_{\text{mix}}+D_{\boldsymbol{\alpha}}$
As neural networks grow, the backward term $(\gamma-1)F$ dominates, so the gap widens in favor of SPSA. In practice, we also amortize $C_{\text{mix}}$ by caching the two blended parameter sets across the paired forward passes.
In this work, GLUE assumes that the architectures are compatible, meaning they share backbones and heads. When the heads differ, we mix the backbone and select or align a single head, optionally refreshing the BatchNorm statistics.
Other limitations include the convex-hull restriction (the target optimum may lie outside the experts’ span), possible misalignment across experts (objectives or preprocessing), hyperparameter sensitivity of SPSA (too large radius $\mu$ of (3) biases the estimate; too small increases noise), and variance growing with $K$ regardless of classical SPSA convergence-to-stationary-point guarantees. These can be mitigated by light calibration, averaging a few directions, or modest schedules, but remain trade-offs.
7 Conclusion
We presented GLUE, a forward-only approach that learns a single deployable prior by mixing pretrained experts in a convex manner and optimizing the mixture via two-point SPSA. The method avoids backpropagation and is accompanied by cost and variance analyses that explain its efficiency and stability. Empirically, GLUE delivers consistent improvements over data-size weighting and proxy-metric baselines while remaining lightweight for deployment when compared to the learned full gradient mixing baseline.